# Next-Gen Interconnection Systems with Compute Express Link: a Comprehensive Survey

### v1

@@ VI Future Research
VI-A CXL Based Memory Extension
VI-A1 Memory Interleaving
@@ VI-A2 Function Offloading
VI-B CXL Based Disaggregation System
VI-B1 CXL Switch
VI-B2 System Model
VI-B3 System Cost
VI-B4 Virtualization
VI-B5 Cross-Rack Interconnect
VI-C CXL Based Distributed Shared Memory
VI-C1 Security
VI-C2 Scalability
VI-C3 Application
VI-C4 Failure Tolerance
VI-D CXL Based Unified Memory
@@ VI-D1 GPU Memory Extension
VI-D2 Heterogeneous Device Collaboration
@@
VI-E CXL Based Independent Memory
CXL memory is located on a separate device, which has independent power supply. This is both a challenge and an opportunity. As for the challenge, we have introduced the new fault tolerance model in section IV-C. As for the opportunity, CXL memory is independent of host, the status of the host or process does not affect the CXL memory, which gives us great research space.
VI-E1 Stateful Workloads Upgrade or Migration
@@
VI-E2 Workloads Cold Start
@@ VI-E3 CXL File System
VII Memory-Centric Computing in CXL Fabric
@@
We define 2 distance, Intra-Cell and Inter-Cell distance. For intra-cell, we can offload function on memory devices to avoid data movement, as same as the existing works in section III-C. For inter-cell, we schedule computing devices according to the location of the data, prioritizing the computing devices within the same cell to reduce the distance.
@@
VIII Conclusion
This paper studies the next-generation interconnection systems based on CXL. We classify the existing research into three categories according to the interconnection types: Memory Expansion and Disaggregation, Distributed Shared Memory, and Unified Memory. Memory Expansion and Disaggregation focuses on processors and memory devices, attempting to address memory wall at both single-machine and cluster scales. Distributed Shared Memory targets cross-node processors, blurring the boundary between single machines and clusters. Unified Memory addresses heterogeneous accelerators, enabling more efficient cooperation in heterogeneous computing systems. Finally, we discuss future research trends in CXL-based interconnection systems and envision a data-centric computing paradigm. We hope this paper provides newcomers with a general understanding and offers fresh inspiration for researchers in the CXL community.

### v2

@@ VIII Future Research
VIII-A CXL Based Memory Extension
VIII-A1 Memory Interleaving
@@ VIII-A2 Function Offloading
VIII-B CXL Based Unified Memory
@@ VIII-B1 GPU Memory Extension
VIII-B2 Heterogeneous Device Collaboration
@@
VIII-C CXL Based Independent Memory
CXL memory is located on a separate device, which has independent power supply. This is both a challenge and an opportunity. As for the challenge, we have introduced the new fault tolerance model in section VII-C. As for the opportunity, CXL memory is independent of host, the status of the host or process does not affect the CXL memory, which gives us great research space.
VIII-C1 Stateful Workloads Upgrade or Migration
@@
VIII-C2 Workloads Cold Start
@@ VIII-C3 CXL File System
VIII-D CXL Based Disaggregated System
VIII-D1 CXL Switch
VIII-D2 System Model
VIII-D3 System Cost
VIII-D4 Virtualization
VIII-D5 Cross-Rack Interconnect
VIII-E CXL Based Distributed Shared Memory
VIII-E1 Security
VIII-E2 Scalability
VIII-E3 Application
VIII-E4 Failure Tolerance
IX Memory-Centric Computing in CXL Fabric
@@
We define 2 distance, Intra-Cell and Inter-Cell distance. For intra-cell, we can offload function on memory devices to avoid data movement, as same as the existing works in section IV. For inter-cell, we schedule computing devices according to the location of the data, prioritizing the computing devices within the same cell to reduce the distance.
@@
X Conclusion
This paper studies the CXL based next-generation computing systems from single-machine to distributed.
In single-machine computing systems, this paper classify the existing research into two categories according to the interconnection types: Memory Expansion and Unified Memory. Memory Expansion focuses on processors and memory devices, attempting to address memory wall with tiered memory system and near-memory processing. Unified Memory focuses on processors and accelerators, enabling more efficient cooperation in heterogeneous computing systems.
In distributed computing systems, this paper discusses how to build high-performance disaggregation systems based on CXL infrastructure, enabling flexible and efficient resource utilization. In addition, this paper also talks about distributed shared CXL memory, which improves the efficiency of communication and synchronization in distributed systems.
Finally, this paper discusses future research trends in CXL-based computing systems and envision a data-centric computing paradigm. We hope this paper provides newcomers with a general understanding and offers fresh inspiration for researchers in the CXL community.